OpenStudy (anonymous):

How do I find the intercepts of 4x^2+y=4

OpenStudy (hba):

@AJQ You have to find x and y intercept Right ?

OpenStudy (hba):

For x intercept y=0 For y intercept x=0 Best of luck :)

OpenStudy (anonymous):

Yes and when I find the X- intercept I need to get rid of the exponent right?

OpenStudy (anonymous):

Do I factor it out?

OpenStudy (hba):

Show me your working :)

OpenStudy (anonymous):

4x^2 + 0=4 4x^2-4=0 (2x+2) (2x-2) 2x+2=0; 2x-2=0 2x/2= -2/2 x=-1 2x/2=2/2 x=1

OpenStudy (hba):

See this 4x^2=4 x^2=1 x=+-1 Got it ?

OpenStudy (anonymous):

No how did you get x^2=1

OpenStudy (anonymous):

divide 4 on both sides?

OpenStudy (hba):

Yes.

OpenStudy (anonymous):

I got it! Thanks!

OpenStudy (hba):

When we remove the sqrt we get +- 1

OpenStudy (hba):

You are welcome :)

OpenStudy (hba):

You can press the best response button if you think my answer was the best :)

OpenStudy (hba):

If you are done with the question please close it then :)